OpenStudy (anonymous):

Will Medal and fan PLEASE HELP Given: F(x) = 3x^2 + 1, G(x) = 2x - 3, H(x) = x F(G(x)) = A: 6x^2-1 B: 12x^2+28 C: 12x^2-36x+28

OpenStudy (kagıtucak):

3.(2x-3)^2+1 just plug 2x-3 instead of x in F(x) =3(4x^2-12x+9)+1=12x^2-36x+28 C

OpenStudy (anonymous):

thank you so much!!!